OpenStudy (anonymous):

Create an expression to find the time for the given distance and rate, and find the value of the expression. d=136 yards; r=17 yards/minutes The time is ................ minutes

OpenStudy (mathmate):

Hint: distance = rate \(\times\) time which makes time = distance \(\div\) rate but check that you are using consistent units, otherwise all quantities must be converted to the same set of units.

OpenStudy (mathmate):

I can give an example: Between two stations 240 miles apart, a train travels at 60 mph. How long would the train take before arriving the destination station? Use the formula above: Time = distance / rate = 240 miles / 60 mph =240 / 60 hours = 4 hours. Your question is very similar to the example.

OpenStudy (anonymous):

So in this problem if i do 240 divided by 60 i would get the answer 4 for the problem that you should

OpenStudy (anonymous):

distance equals rate times time (d=rt) plug in the values and solve

OpenStudy (anonymous):

so if i do 136 divided by 17 i should get my answer

OpenStudy (anonymous):

Thanks for your help @mathmate

OpenStudy (mathmate):

Yes, the division will give the right numerical value. But it is important to check the units. Since they are both in yards, and the time unit is in minutes, then they are in consistent units. The result will be in minutes. For the units, \(\large yards\div\frac{yards}{minutes}\) =\(\large yards \times \frac{minutes}{yards}\) =\(\large minutes\)

OpenStudy (mathmate):

You're welcome! :)